Please evaluate x/square root 1+x^2 over the interval [0,1]
I believe this is a definite interger

Answer by stanbon(48558)

Please evaluate x/square root 1+x^2 over the interval [0,1]
I believe this is a definite integral
----------------
Integral[(x /sqrt(1+x^2))dx] over [0,1]
=(1/2)Integral [(1+x^2)^(-1/2)(2x)dx over [0.1]
=(1+x^2)^(1/2) over [0,1]
=sqrt[1+1}-sqrt[1+0]
=sqrt2 -1
=0.414...
Cheers,
Stan H.